OpenStudy (anonymous):

Find the Laplace Transform using trig identities for: t sin(2t) sin(5t) this is what I've got so far: (1/4i)[(s-7i)^-2 + (s+7i)^-2 - (s-3i)^-2 - (s+3i)^-2]

OpenStudy (anonymous):

\[2sinAsinB = \cos(A-B) - \cos(A+B)\] \[tsin(2t)\sin5(t) = \frac{ 1 }{ 2 }t(\cos(-3t) - \cos(7t))\] then take laplace of that. is this what you did? not sure what your step is though and why you have i

OpenStudy (anonymous):

this table is very good btw, in case you don't have a good one already: http://tutorial.math.lamar.edu/pdf/Laplace_Table.pdf you'll want to refer to number 10 or number 30 [the general case]

OpenStudy (anonymous):

thank you i'll try to solve it using what you've taught me

OpenStudy (anonymous):

I used the formula that says: \[\sin(a)=(e^{ai}-e^{-ai})/2i\]

OpenStudy (anonymous):

Now by using \[(1/2)t [\cos(-3t) - \cos(7t)]\] I got \[-20s/[(s^2+9)(s^2+49)]\] is it the right answer ??

OpenStudy (anonymous):

i believe that is sinh(a) not sin(a) \[\frac{ 1 }{ 2 }t[\cos(-3t) - \cos(7t)]\ using\ equation\ 10\] \[\frac{ 1 }{ 2 }\left( \frac{ s^2 - 9 }{ (s^2 +9)^2 }-\frac{ s^2 - 49 }{ (s^2 + 49)^2 } \right)\]

OpenStudy (anonymous):

thank you now i can understand what you were doing to solve it but the only problem is i didn't get how you compute the function from the above trig identity \[2sinAsinB=\cos(A−B)−\cos(A+B)\] into \[tsin(2t)\sin(5t)=(1/2)t[\cos(−3t)−\cos(7t)]\]

OpenStudy (anonymous):

A = 2t B = 5t cos(A-B) = cos(2t - 5t) = cos(-3t) cos(A+B) = cos(2t + 5t) = cos(7t)

OpenStudy (anonymous):

thank you so much, that was very helpful =)